# 22email: wai

### v2

@@ 3 Experimental setup
Baselines. We use non-GR and GR baselines. Following [lee2023glen], we use the following non-GR baselines: BM25 [robertson2009probabilistic], DPR [karpukhin2020dense], SentenceT5 [ni2022sentence], and GTR-base [ni2022large]. We use the following GR baselines (training-based indexing): (i) SEAL [bevilacqua2022autoregressive] learns to generate n-grams-based docids and applies FM-index [ferragina2000opportunistic]. (ii) DSI [tay2022transformer] learns to generate numeric identifiers. (iii) DSI-QG [zhuang2022bridging] augments DSI training by using pseudo queries; we replicate DSI-QG using the pseudo query generator provided by the original paper. (iv) DSI-QG (InPars) uses the pseudo query generator from InPars [bonifacio2022inpars]. (v) TOME [ren2023tome] learns to generate document URLs. (vi) GLEN [lee2023glen] learns dynamic lexical docids. (vii) GenRET [sun2024learning] learns to assign numeric docids based on an auto-encoding scheme. (viii) NOVO [wang2023novo] learns interpretable docids.

### v3

@@ 2 Methodology
Table 4 in the appendix gives an example of 10 distinct docids generated by Few-Shot GR for a specific document in NQ320K [lee2023glen, sun2024learning, tay2022transformer].
@@ 3 Experimental setup
Baselines. We use non-GR and GR baselines. Following [lee2023glen], we use the following non-GR baselines: BM25, DPR [karpukhin2020dense], SentenceT5 [ni2022sentence], and GTR-base [ni2022large]. We use the following GR baselines (training-based indexing): (i) SEAL [bevilacqua2022autoregressive] learns to generate n-grams-based docids and applies FM-index [ferragina2000opportunistic]. (ii) DSI [tay2022transformer] learns to generate numeric identifiers. (iii) DSI-QG [zhuang2022bridging] augments DSI training by using pseudo queries; we replicate DSI-QG using the pseudo query generator provided by the original paper. (iv) DSI-QG (InPars) uses the pseudo query generator from InPars [bonifacio2022inpars]. (v) TOME [ren2023tome] learns to generate document URLs. (vi) GLEN [lee2023glen] learns dynamic lexical docids. (vii) GenRET [sun2024learning] learns to assign numeric docids based on an auto-encoding scheme. (viii) NOVO [wang2023novo] learns interpretable docids.
@@ 5 Conclusions & Future Work
Appendix
Case study of docids generated by Few-Shot GR. Table 4 gives an example of 10 distinct docids generated by Few-Shot GR for a specific document in NQ320K. It shows that docids generated by Few-Shot GR are various.
Table 4: Pseudo queries and corresponding docids generated by Few-Shot GR using Llama-3-8B-Instruct for a document from NQ320K.
[…]
In accounting, minority interest (or non-controlling interest) is the portion of a subsidiary corporation’s stock that is not owned by the parent corporation. The magnitude of the minority interest in the subsidiary company is generally less than 50% of outstanding shares, or the corporation would generally cease to be a subsidiary of the parent. What is minority interest in accounting? minority-interest-accounting
What is non-controlling interest in accounting? non-controlling-interest-accounting
How is minority interest defined in accounting? minority-interest-definition
How is minority interest calculated in accounting? minority-interest-calculation
What is the significance of minority interest in accounting? minority-interest-significance
How does minority interest affect financial statements in accounting? minority-interest-financial-statements
How is minority interest treated in consolidated financial statements in accounting? minority-interest-consolidated-financial-statements
What is the impact of minority interest on the parent company’s earnings per share in accounting? minority-interest-impact-eps
How is minority interest reported in financial statements of a subsidiary company in accounting? minority-interest-reporting-subsidiary
What is the role of minority interest in a parent corporation’s financial statements? minority-interest-parent-corporation